with what speed must a ball be thrown down for it to bounce 10 m higher than its original level? assume that no enegry loss in striking the ground.

Grade:11

1 Answers

Khimraj
6 years ago
Let intial speed is v and initial height of bounce is h
½mv= mgh    – (1)
h = v2/2g
Let required velocity be V, then
½mV= mg(h+10)  – (2)
dividing equation (2) by (1)
V2/v= 1 + 10/h
V= v+ (v2)(10*2g/v2)
V2 = v2 + 200
V = (v2 + 200)1/2 
Hope it clears.